### v3

@@ 6.1 Instance Set
We use the instance generator developed by Dall’Olio and Kolisch (2023) to construct a set of test instances of different complexity. A predefined number of flights and their characteristics, such as time windows and parking positions, are generated based on realistic assumptions from Munich Airport. In the following, we briefly elaborate on parameters that have been varied throughout the generation.
In order to generate instances of different sizes, we vary the length of the planning horizon between 60, 90, and 120 minutes and generate 10, 20, and 30 flights per hour. The underlying time grid consists of equidistant time steps with a length of 2 minutes. Each type of airplane served at Munich Airport can be loaded using one of up to 3 different team formations, called slow, intermediate, and fast mode. Faster modes use up more workers but require less time to (un)load an aircraft. Not every plane can be loaded with all 3 modes, especially smaller planes typically only support slow or fast modes. We note that the same properties hold for unloading tasks.
@@
All further properties, such as flight schedules and task execution times, are calculated as described by Dall’Olio and Kolisch (2023). For a precise description of the process, we refer to Section 6.1 and Appendices E and F of their paper.
We approximate the distribution of travel times between locations using a quadratic regression model that aims to predict travel times based on several exogenous factors. For this purpose, we collected real-world data on travel times of various ground operating vehicles, such as passenger buses and baggage handling vehicles, at Terminal 2 of Munich Airport for the entire year of 2024. These values, if needed, are adjusted to baggage handling vehicles by scaling them proportionally by their respective average speeds. We then categorized all datapoints based on several features, which in previous studies were identified to have a significant impact on travel times: weather conditions (cloudy, sunny, rainy, snowy), time during the day (on-peak or off-peak) and the week (weekday or weekend, holiday season yes or no), number of required apron crossings to get from origin to destination gate, distance, and best-case travel time. To estimate travel time distributions for a given gate pair under specified external conditions, we then apply the quadratic regression model to predict the percentual deviation from the shortest travel time. To account for variability, we then identify historical observations from other gate pairs with similar deterministic travel times, distances, and numbers of crossings. For these similar observations, we calculate the prediction errors of the percentual delay under their respective external conditions. Finally, we construct the travel time distribution for the target gate pair as follows. For each residual obtained from the aforementioned similar observations, we add the residual to the predicted percentual delay for the target gate pair and multiply the result by the best-case travel time. The resulting set of datapoints is then used to derive an empirical distribution.
@@ 6.2 Influence of the Algorithm’s Features
From a methodological point of view, our solution approach differs from the ones typically used for variants of vehicle routing or team formation problems in three ways: dynamically switching between two master problem formulations, branching on task finish times and an exact separation of rank-1 CGCs. In the following, these are called the algorithm’s features. In this section, we compare a total of six solver configurations: we first enable and then disable all three features, resulting in what are called ‘full’ and ‘basic’ configurations (or algorithms), respectively. Furthermore, we implemented the branch-and-price approach proposed by Yuan, Liu, and Jiang (2015), which differs from our approach in several ways. First, they do not use cutting planes and employ branching on arcs as their sole branching rule. Second, they apply a greedy and a neighborhood search heuristic to find negative columns before applying a labeling algorithm. Third, they only use a disaggregated master problem formulation, i.e., they do not switch between master problem formulations, and always return all negative columns found rather than just the most negative column. Fourth, while we aggregate multiple pricing networks into one by comparing offset reduced costs and adding an additional restriction (24) on the depot leave times, they demand reduced cost dominance for all skill compositions. Thus, unlike us, they implicitly solve all pricing networks exactly.
Moreover, we disable exactly one feature to obtain three additional variants of our solution method called ‘no DRMP’, ‘no CGCs’ and ‘no branching on task finish times’ (also abbreviated as ‘no branching’). For instance, the configuration ‘no DRMP’ uses branching on task finish times and adds up to 12 CGCs at the root node, but does not dynamically switch to the DRMP master problem formulation. We note that we do not include solver configurations with exactly two features enabled in our analysis, as the results do not fundamentally differ from the ones presented in the following.
@@
We configure the aforementioned features as follows. When branching, we first try to branch on task finish times by using the procedure described in Section 5.5. If that is not possible, we look for branches on vehicle counts and, if this also fails, we use the fallback option of branching on variables. Furthermore, whenever a disaggregated-infeasible solution is found, we switch to solving the DRMP formulation at the current node and its sibling node. Additionally, child nodes of DRMP nodes inherit their master problem’s type, i.e., they are also solved using the DRMP formulation. Finally, we only search for violated CGCs at the root node and add up to 12 cuts. The MIP required to identify such cuts, as described by Fischetti and Lodi (2007) is solved using Gurobi with a time limit of 0.3 seconds. For each instance, a hard time limit of 180 seconds is imposed. If no optimal solution has been found, an upper bound is obtained by the procedure described in Section 5.7. We note that preliminary studies have shown that adding up to 12 cuts, on average, provides an optimal trade-off between lower bound improvements and an increase in computational complexity.
@@
Each of the aforementioned 5 configurations is used to solve the instance set generated as described in Section 6.1, amounting to a total of 1,215 instances per configuration. In total, 615 instances are infeasible. We omit said instances and compare the ascribed feature configurations on the remaining 600 instances. Appendix 15 contains additional details on the characteristics of all feasible instances.
In order to distinguish between easy and hard instances, we split the set of test instances into three categories. In preliminary studies, the worker strength has proven to have a significant impact on an instance’s complexity. Therefore, we consider instances with a worker strength between 0.3 and 0.5 as ‘hard’, while instances with a strength of 0.6 or 0.7 are seen as ‘medium’ and 0.8 or 0.9 as ‘easy’. In fact, almost all instances in the latter category were solved within a few seconds, whereas instances in the hard category are frequently not solved to optimality within the set time limit. We note that all instances with a worker strength of 0.2 or less are infeasible, while only 2 instances with a worker strength of 0.3 are feasible.
Unless otherwise stated, all tables in this section, contain average values. The number of instances on which the following analyses are based can be seen in Appendix 16.
@@
Table 4: Comparison of Yuan, Liu, and Jiang (2015) with basic and full solver configuration
@@
Easy Full 100.00% 0.00% 2.86 2.86 1.53s
Basic 82.76% 6.63% 2.86 2.12 33.30s
Yuan et al. 92.34% 2.12% 2.86 2.62 20.28s
Medium Full 90.43% 0.29% 15.54 15.48 31.84s
Basic 57.89% 7.48% 15.59 14.29 84.80s
Yuan et al. 63.46% 5.25% 15.54 14.56 80.45s
Hard Full 64.52% 2.77% 67.62 64.60 84.99s
Basic 49.61% 4.29% 69.19 64.84 106.08s
Yuan et al. 42.40% 4.95% 68.92 64.66 122.18s
All Full 89.23% 0.68% 20.84 20.19 29.61s
Basic 67.06% 6.44% 21.27 19.57 66.68s
Yuan et al. 71.72% 3.81% 21.20 19.86 62.79s
Column “% Opt” contains the percentage of instances that have been solved to optimality. It can be seen that the full configuration solves 22% more instances to optimality than the basic configuration. Furthermore, it returns significantly better average optimality gaps for all instance classes, where this effect decreases with increasing worker strength. While the full configuration returns substantially better lower bounds for all instance classes, the upper bounds also increased slightly for harder instances. Moreover, runtimes for small and medium instances reduce drastically, while larger instances are solved around 22 seconds faster on average. Additionally, it can be observed that, while the algorithm by Yuan, Liu, and Jiang (2015) is able to outperform the basic configuration for easy instances, it struggles to do so for medium and hard instances. Finally, the proposed full configuration manages to return the best runtimes, bounds, and optimality percentages out of all compared approaches for all complexity classes.
@@
Figure 3 visualizes the frequencies of non-zero optimality gap percentages for all five considered solver configurations in a box plot. The red line is the median, while the boxes are limited by the 25%- and 75%-quantiles. The lower and upper whiskers are calculated as $q_{0.25}-1.5\cdot IQR$ and $q_{0.75}+1.5\cdot IQR$, respectively, where $IQR=q_{0.75}-q_{0.25}$ is the interquartile range. Clearly, the basic configuration has a high median optimality gap, as well as a comparable large set of outliers, sometimes even terminating with a gap of close to 70%. These metrics generally improve when enabling two of the three algorithm’s features, while the best results can be observed when using the full solver configuration. Additionally, larger gap percentages occur rather frequently when using the basic configuration, as all instances lie within their respective whiskers. Hence, the full solver configuration returns the best and most stable results from all compared strategies.
@@
Easy Full 100.00% 0.00% 2.86 2.86 1.39s 1.28 1.54 0.99s 0.041s
No Branching 99.23% 0.27% 2.86 2.85 1.47s 1.93 6.50 0.45s 0.039s
Medium Full 90.43% 0.29% 15.54 15.48 5.79s 11.39 36.73 0.87s 0.111s
No Branching 73.68% 2.57% 15.58 15.04 7.00s 7.59 49.43 1.14s 0.219s
Hard Full 64.52% 2.77% 67.62 64.60 16.17s 21.00 65.24 1.30s 0.242s
No Branching 56.00% 3.73% 69.30 65.29 17.95s 17.95 50.32 1.90s 0.386s
All Full 89.23% 0.68% 20.84 20.19 4.74s 7.01 27.22 1.09s 0.108s
No Branching 81.18% 1.81% 21.29 20.25 5.31s 5.83 30.78 1.34s 0.175s
Table 5 compares the results obtained by the full solver configuration with the configuration ‘no branching’, i.e., branching on task finish times is disabled. Columns “Ropt” and “Nopt” describe the runtime and the number of explored nodes until optimality has been proven. We note that, unlike all other values in the above table, the baseline for these two metrics is not the entire instance set of a fixed instance class, but instead the set of instances that has been solved to optimality by both solver configurations. Column “Nexp” describes the number of explored nodes, while “Rnode” and “Rpr” contain the runtime per explored node and the runtime per pricing iteration, respectively. The average solution quality improves when the branching strategy is enabled, as 8.05% additional instances are solved to optimality, and the average gap is reduced by around 1.27%. Furthermore, solving a single node requires, on average, 25% less time, where this improvement increases for harder instances. This can be explained by significant time savings in the pricing step, where the runtime decreases from 0.386 seconds per iteration to 0.242 seconds.
For easy and medium instances, the full solver configuration requires less time to prove optimality and explores around 40% to 50% fewer nodes to do so, respectively. For the set of hard test instances, using branching on task finish times improves the solution quality noticeably, while optimality gaps are reduced by 1.04%. However, around 20% more nodes are required to prove optimality. This indicates that for larger instances, branching on task finish times is able to provide very good lower bounds, but does not excel at proving optimality. In summary, task finish times appear to be an efficient basis for branching decisions, as they significantly simplify the pricing step. For more complex instances, further studies on a potential tail-off effect on the quality of lower bounds and possible ways to mitigate this issue are needed.
@@
Easy Full 100.00% 0.00% 2.86 2.86 1.44s 1.49 1.00 4.70 1.00
No DMP 100.00% 0.00% 2.86 2.86 1.50s 1.85 3.00 0.00 1.00
Medium Full 90.43% 0.29% 15.54 15.48 8.59s 18.63 44.50 7.90 1.12
No DMP 86.12% 0.57% 15.55 15.46 8.21s 17.57 88.80 0.00 19.44
Hard Full 64.52% 2.77% 67.62 64.60 24.18s 38.84 123.40 6.10 2.09
No DMP 63.71% 2.91% 67.55 64.42 20.14s 35.33 131.60 0.00 9.36
All Full 89.23% 0.68% 20.84 20.19 7.05s 12.48 63.40 7.20 1.38
No DMP 87.54% 0.81% 20.83 20.14 6.38s 11.82 94.30 0.00 15.18
@@
Easy Full 100.00% 0.00% 2.86 2.86 0.99 0.71 92.72% 2.84
No CGCs 96.55% 0.46% 2.86 2.82 0.44 0.46 47.89% 1.97
Medium Full 90.43% 0.29% 15.54 15.48 0.87 0.76 50.72% 14.79
No CGCs 83.17% 1.22% 15.28 15.13 0.44 0.43 13.44% 13.13
Hard Full 64.52% 2.77% 67.62 64.60 1.30 1.00 29.03% 62.88
No CGCs 60.80% 2.85% 68.83 65.66 0.68 0.62 7.20% 62.26
All Full 89.23% 0.68% 20.84 20.19 1.09 0.88 64.65% 19.58
No CGCs 84.34% 1.23% 21.09 20.36 0.54 0.51 27.44% 18.56
Table 7 analyzes the results obtained by the full configuration and compares it to the ‘no CGCs’ configuration, i.e., no CGCs are added. Column “rootLB” contains the lower bound obtained after solving the root node, while column “RneR” depicts the algorithm’s runtime per node, excluding the root node. On average over all test instances, the usage of Gomory cuts improves the final lower bounds, allowing us to solve 4% more instances to optimality and closing the optimality gap by an additional 0.55%. Furthermore, the percentage of instances solved at the root node more than doubles from 27.44% to 64.65%, while it even increases four-fold for hard instances. Moreover, the lower bound at the root node improves by around 6%. Additionally, column “RnerR” shows that CGCs significantly increase the runtime per non-root node, doubling from 0.54 seconds to 1.09 seconds on average. Altogether, Gomory cuts aid to provide excellent bounds early on during the solving procedure. However, the increase in computational complexity, especially during the pricing step, implies a trade-off between bound quality and computational complexity. Especially for larger instances, carefully separating and electing CGCs plays a crucial role in the approach’s efficiency.
To conclude this section, we showed that our developed solution strategy is able to significantly improve the ‘basic’ configuration first described by Dall’Olio and Kolisch (2023), both with respect to optimality guarantees and bound quality. While the impact of every single configuration’s component seems to be rather small, combining our three core features and using them simultaneously greatly benefits the resulting algorithm’s performance.

### v4

@@ 6.1 Instance Set
We use the instance generator developed by dallolio_temp to construct a set of test instances of different complexity. In short, a flight plan for a given time horizon is generated based on real flight schedules and aircraft usage. The workforce size depends on the instance and varies between 10% and 90% of the workforce required to trivially solve the instance (i.e., each task is started at its earliest possible time and is executed using the profile that requires the shortest task execution time). Profiles, skill compositions, and task execution times are defined based on aircraft manual data. For a detailed summary of the process of test instance generation, we refer to Section 6.1 of dallolio_temp. In the following, we briefly elaborate on parameters that have been varied throughout the generation.
In order to generate instances of different sizes, we vary the length of the planning horizon, as well as the frequency of flights. At Munich Airport, shift planning is typically done in three shifts (early, late, and night). Furthermore, the frequency of arrivals and departure can be categorized into low-frequency (i.e. off-peak) and high-frequency (i.e. peak) time periods. Each shift contains up to two such peaks, and each peak lasts for 1 to 2 hours and contains between 10 and 30 flights per hour. Because shift planning is done based on the workforce demand in peak times, the problem at hand becomes trivial to solve outside of peak periods due to the abundance of workforce. Hence, in the following, we focus on the aforementioned peak periods and create instances with a planning horizon between 60, 90, and 120 minutes and 10, 20, and 30 flights per hour. The underlying time grid consists of equidistant time steps with a length of 2 minutes. Each type of airplane served at Munich Airport can be loaded using one of up to 3 different team formations, called slow, intermediate, and fast mode. Faster modes use up more workers but require less time to (un)load an aircraft. Not every plane can be loaded with all 3 modes, especially smaller planes typically only support slow or fast modes. We note that the same properties hold for unloading tasks.
@@
All further properties, such as flight schedules and task execution times, are calculated as described by dallolio_temp. For a precise description of the process, we refer to Section 6.1 and Appendices E and F of their paper.
We approximate the distribution of travel times between locations using a quadratic regression model that aims to predict travel times based on several exogenous factors. For this purpose, we collected real-world data on travel times of various ground operating vehicles, such as passenger buses and baggage handling vehicles, at Terminal 2 of Munich Airport for the entire year of 2024. These values, if needed, are adjusted to baggage handling vehicles by scaling them proportionally by their respective average speeds. We then categorized all datapoints based on several features, which in preliminary internal studies were identified to have a significant impact on travel times: weather conditions (cloudy, sunny, rainy, snowy), time during the day (on-peak or off-peak) and the week (weekday or weekend, holiday season yes or no), number of required apron crossings to get from origin to destination gate, distance, and best-case travel time. To estimate travel time distributions for a given gate pair under specified external conditions and a given time bin, we then apply the quadratic regression model to predict the percentual deviation from the shortest travel time. To account for variability, we then identify historical observations from other gate pairs in the same time bin with similar deterministic travel times, distances, and numbers of crossings. For these similar observations, we calculate the prediction errors of the percentual delay under their respective external conditions. Finally, we construct the travel time distribution for the target gate pair as follows. For each residual obtained from the aforementioned similar observations, we add the residual to the predicted percentual delay for the target gate pair and multiply the result by the best-case travel time. The resulting set of datapoints is then used to derive an empirical distribution for the given edge in the give time bin.
@@ 6.2 Influence of the Algorithm’s Features
From a methodological point of view, our solution approach differs from the ones typically used for variants of vehicle routing or team formation problems in three ways: dynamically switching between two master problem formulations, branching on task finish times and an exact separation of rank-1 CGCs. In the following, these are called the algorithm’s features. In this section, we compare a total of six solver configurations: we first enable and then disable all three features, resulting in what are called ‘full’ and ‘basic’ configurations (or algorithms), respectively. Furthermore, we implemented the branch-and-price approach proposed by yuan2015branch, which differs from our approach in several ways. First, they do not use cutting planes and employ branching on arcs as their sole branching rule. Second, they apply a greedy and a neighborhood search heuristic to find negative columns before applying a labeling algorithm. Third, they only use a disaggregated master problem formulation, i.e., they do not switch between master problem formulations, and always return all negative columns found rather than just the most negative column. Fourth, while we aggregate multiple pricing networks into one by comparing offset reduced costs and adding an additional restriction (26) on the depot leave times, they demand reduced cost dominance for all skill compositions. Thus, unlike us, they implicitly solve all pricing networks exactly. While yuan2015branch consider a simplified variant of our problem and assume stochastic service times instead of travel times, their proposed approach can be adapted to solve the problem at hand.
Moreover, we disable exactly one feature to obtain three additional variants of our solution method called ‘no DRMP’, ‘no CGCs’ and ‘no branching on task finish times’ (also abbreviated as ‘no branching’). For instance, the configuration ‘no DRMP’ uses branching on task finish times and adds up to 12 CGCs at the root node, but does not dynamically switch to the DRMP master problem formulation. We note that we do not include solver configurations with exactly two features enabled in our analysis, as the results do not fundamentally differ from the ones presented in the following. Furthermore, all results reported in this subsection are based on a test instance set with $\gamma=0.9$. While we also conducted analogous analyses for the same test instances using $\gamma=0.5$, $\gamma=0.7$ and $\gamma=1$, the performance of the proposed algorithm does not fundamentally differ.
@@
We configure the aforementioned features as follows. When branching, we first try to branch on task finish times by using the procedure described in Section 5.5. If that is not possible, we look for branches on vehicle counts and, if this also fails, we use the fallback option of branching on variables. Furthermore, whenever a disaggregated-infeasible solution is found, we switch to solving the DRMP formulation at the current node and its sibling node. Additionally, child nodes of DRMP nodes inherit their master problem’s type, i.e., they are also solved using the DRMP formulation. Finally, we only search for violated CGCs at the root node and add up to 12 cuts. The MIP required to identify such cuts, as described by fischetti2007optimizing is solved using Gurobi with a time limit of 0.3 seconds. For each instance, a hard time limit of 180 seconds is imposed. If no optimal solution has been found, an upper bound is obtained by the procedure described in Section 5.7. We note that preliminary studies have shown that adding up to 12 cuts and imposing a runtime limit of 0.3 seconds, on average, provides an optimal trade-off between lower bound improvements and an increase in computational complexity.
@@
Each of the aforementioned 5 configurations is used to solve the instance set generated as described in Section 6.1, amounting to a total of 1,215 instances per configuration. In total, 614 instances are infeasible. We omit said instances and compare the ascribed feature configurations on the remaining 601 instances. H contains additional details on the characteristics of all feasible instances.
In order to distinguish between easy and hard instances, we split the set of test instances into three categories. In preliminary studies, the worker strength has proven to have a significant impact on an instance’s complexity. Therefore, we consider instances with a worker strength between 0.3 and 0.5 as ‘hard’, while instances with a strength of 0.6 or 0.7 are seen as ‘medium’ and 0.8 or 0.9 as ‘easy’. In fact, almost all instances in the latter category were solved within a few seconds, whereas instances in the hard category are frequently not solved to optimality within the set time limit. We note that all instances with a worker strength of 0.2 or less are infeasible, while only 3 instances with a worker strength of 0.3 are feasible.
Unless otherwise stated, all tables in this section, contain average values. The number of instances on which the following analyses are based can be seen in I.
@@
Table 2: Comparison of yuan2015branch with basic and full solver configuration
@@
Easy Full 100.00% 0.00% 2.77 2.77 2.29
Basic 82.38% 6.48% 2.77 2.06 33.47
Yuan et al. 94.25% 1.15% 2.77 2.60 18.07
Medium Full 91.08% 0.35% 14.95 14.91 26.21
Basic 62.44% 6.48% 14.96 13.83 79.67
Yuan et al. 68.54% 3.66% 14.95 14.25 71.34
Hard Full 70.08% 1.69% 67.91 65.95 79.24
Basic 47.24% 4.10% 68.20 64.52 111.22
Yuan et al. 42.52% 3.91% 67.85 64.53 121.34
All Full 90.52% 0.48% 20.85 20.42 27.03
Basic 67.89% 5.98% 20.92 19.43 66.27
Yuan et al. 74.21% 2.62% 20.84 19.81 58.77
Column “% Opt” contains the percentage of instances that have been solved to optimality. It can be seen that the full configuration solves 23% more instances to optimality than the basic configuration. Furthermore, it returns significantly better average optimality gaps for all instance classes, where this effect decreases with increasing worker strength. While the full configuration returns substantially better lower bounds for all instance classes, the upper bounds also increased slightly for harder instances. Moreover, runtimes for small and medium instances reduce drastically, while harder instances are solved around 22 seconds faster on average. Additionally, it can be observed that, while the algorithm by yuan2015branch is able to outperform the basic configuration for easy and medium instances, it struggles to do so for hard instances. Finally, the proposed full configuration manages to return the best runtimes, bounds, and optimality percentages out of all compared approaches for all complexity classes.
@@
Figure 4 visualizes the frequencies of non-zero optimality gap percentages for all five considered solver configurations in a box plot. The red line is the median, while the boxes are limited by the 25%- and 75%-quantiles. The lower and upper whiskers are calculated as $q_{0.25}-1.5\cdot IQR$ and $q_{0.75}+1.5\cdot IQR$, respectively, where $IQR=q_{0.75}-q_{0.25}$ is the interquartile range. Clearly, the basic configuration has a high median optimality gap, as well as a comparable large set of outliers, sometimes even terminating with a gap of close to 70%. These metrics generally improve when enabling two of the three algorithm’s features, while the best results can be observed when using the full solver configuration. Additionally, larger gap percentages occur rather frequently when using the basic configuration, as all instances lie within their respective whiskers. Thus, the full solver configuration is far superior to the basic configuration as well as to the configuration ’no branching’. Moreover, for configurations ’No CGCs’ and ’No DMP’, although the improvements dimish, the full configuration still yields fewer outliers, a smaller median gap, shorter whiskers, and a comparable IQR. However, Figure 4 only analyzes the distribution of optimality gaps, neglecting differences in other metrics that influence the performance of an algorithm. Hence, in the following, we evaluate additional metrics, including runtime, average upper and lower bounds, and the percentage of instances solved to optimality, to further assess the benefits of the proposed approach’s features.
@@
Easy Full 100.0% 0.00% 2.77 2.77 2.28 1.63 1.64 0.69s 0.048s
No Branching 99.23% 0.05% 2.77 2.77 2.65 3.78 5.86 1.40s 0.047s
Medium Full 91.08% 0.35% 14.95 14.91 7.26 7.32 24.01 0.94s 0.116s
No Branching 79.34% 1.96% 14.96 14.56 9.97 12.01 48.73 1.09s 0.184s
Hard Full 70.08% 1.69% 67.91 65.95 22.21 26.59 47.81 1.66s 0.226s
No Branching 58.27% 3.08% 68.21 65.17 23.85 23.86 45.08 2.06s 0.379s
All Full 90.52% 0.48% 20.85 20.42 6.80 7.11 19.33 1.28s 0.110s
No Branching 83.53% 1.37% 20.92 20.13 8.14 9.41 29.34 1.40s 0.165s
Table 3 compares the results obtained by the full solver configuration with the configuration ‘no branching’, i.e., branching on task finish times is disabled. Columns “Ropt” and “Nopt” describe the runtime and the number of explored nodes until optimality has been proven. We note that, unlike all other values in the above table, the baseline for these two metrics is not the entire instance set of a fixed instance class, but instead the set of instances that has been solved to optimality by both solver configurations. Column “Nexp” describes the number of explored nodes, while “Rnode” and “Rpr” contain the runtime per explored node and the runtime per pricing iteration, respectively. The average solution quality improves when the branching strategy is enabled, as 7% additional instances are solved to optimality, and the average gap is reduced by around 0.9%. Furthermore, solving a single node requires, on average, 10% less time, where this improvement increases for harder instances. This can be explained by significant time savings in the pricing step, where the runtime decreases from 0.379 seconds per iteration to 0.226 seconds.
For easy and medium instances, the full solver configuration requires less time to prove optimality and explores around 40% to 60% fewer nodes to do so, respectively. For the set of hard test instances, using branching on task finish times improves the solution quality noticeably, while optimality gaps are reduced by 1.39%. However, around 12% more nodes are required to prove optimality. This indicates that for larger instances, branching on task finish times is able to provide very good lower bounds, but does not excel at proving optimality. In summary, task finish times appear to be an efficient basis for branching decisions, as they significantly simplify the pricing step. For more complex instances, further studies on a potential tail-off effect on the quality of lower bounds and possible ways to mitigate this issue are needed.
@@
Easy Full 100.00% 0.00% 13.45 13.45 48.30s 18.40 1.00 17.4 1.00
No DMP 100.00% 0.00% 13.45 13.45 9.24s 2.80 2.80 0.00 1.00
Medium Full 78.26% 0.21% 19.42 19.34 37.97s 39.47 43.90 7.20 1.09
No DMP 69.57% 1.11% 19.42 19.31 62.78s 43.47 55.00 0.00 8.09
Hard Full 93.33% 0.03% 54.90 54.85 72.26s 70.45 43.00 15.70 1.07
No DMP 73.33% 0.17% 54.91 54.78 62.03s 68.00 73.80 0.00 13.07
All Full 86.05% 0.12% 31.11 31.04 51.80s 47.06 38.60 11.30 1.07
No DMP 74.42% 0.65% 31.11 31.00 53.88s 45.61 55.50 0.00 9.00
@@
Easy Full 100.00% 0.00% 2.77 2.77 1.40s 1.89s 92.34% 2.75
No CGCs 96.55% 0.16% 2.77 2.75 0.53s 0.55s 46.74% 1.90
Medium Full 91.08% 0.35% 14.95 14.91 1.09s 0.95s 56.81% 14.20
No CGCs 85.92% 0.70% 14.95 14.86 0.60s 0.60s 15.49% 12.87
Hard Full 70.08% 1.69% 67.91 65.95 1.66s 1.27s 25.98% 63.18
No CGCs 66.14% 1.46% 67.89 66.07 1.05s 0.96s 4.72% 61.05
All Full 90.52% 0.48% 20.85 20.42 1.40s 1.14s 65.72% 19.58
No CGCs 86.36% 0.62% 20.85 20.42 0.75s 0.72s 26.79% 18.29
Table 5 analyzes the results obtained by the full configuration and compares it to the ‘no CGCs’ configuration, i.e., no CGCs are added. Column “rootLB” contains the lower bound obtained after solving the root node, while column “RneR” depicts the algorithm’s runtime per node, excluding the root node. On average over all test instances, the usage of Gomory cuts improves the final lower bounds, allowing us to solve 4.16% more instances to optimality and closing the optimality gap by an additional 0.14%. Furthermore, the percentage of instances solved at the root node more than doubles from 26.79% to 65.72%, while it even increases six-fold for hard instances. Moreover, the lower bound at the root node improves by around 7%. Additionally, column “RnerR” shows that CGCs significantly increase the runtime per non-root node, doubling from 0.75 seconds to 1.4 seconds on average. Altogether, Gomory cuts aid to provide excellent bounds early on during the solving procedure. However, the increase in computational complexity, especially during the pricing step, implies a trade-off between bound quality and computational complexity. Especially for larger instances, carefully separating and selecting CGCs plays a crucial role in the approach’s efficiency.
To conclude this section, we showed that our developed solution strategy is able to significantly improve the ‘basic’ configuration first described by dallolio_temp, both with respect to optimality guarantees and bound quality. While the impact of every single configuration’s component seems to be rather small, combining our three core features and using them simultaneously greatly benefits the resulting algorithm’s performance.